Find the projection of the vector v onto the subspace S.

S = span (0,1,1), (1,1,0),    v = (3,4,2)


proj_s v = ______


1
Expert's answer
2020-10-27T17:56:24-0400

Given that ,

"S=Span \{ (0,1,1),(1,1,0)\}" and "v=(3,4,2)"

let "u_1=(0,1,1) \ and \ u_2=(1,1,0)"

Then "S=Span\{u_1,u_2\}" .

Since "u_1" and "u_2" are not orthogonal ,first apply the Gram-Schmidt algorithm to find an orthogonal basis for "S" .

Set "w_1=u_1=(0,1,1)"

Then find "w_2=u_2-\frac{<u_2,w_1>}{<w_1,w_1>}w_1" "=(1,1,0)-\frac{1}{2}(0,1,1)"

"=(1,\frac{1}{2},-\frac{1}{2})"

Where "w_1\ and \ w_2" are orthogonal basis of "U" .

Now we have to calculate the Fourier coefficient of "v" with respect to "u_i" i,e,


"c_1=\frac{<v,w_1>}{<w_1,w_1>}=\frac{6}{2}=3"


"c_2=\frac{<v,w_2>}{<w_2,w_2>}=\frac{3+2-1}{1+\frac{1}{4}+\frac{1}{4}}=\frac{4}{(\frac{3}{2})}" "=\frac{8}{3}"


Then projection "Proj(v,U)" "=c_1w_1+c_2w_2"

"=3(0,1,1)+\frac{8}{3}(1,\frac{1}{2},-\frac{1}{2})"

"=(0,3,3)+(\frac{8}{3},\frac{4}{3},-\frac{4}{3})"

"=(\frac{8}{3},\frac{13}{3},\frac{5}{3})"